# Toward a Physical Theory of Intelligence

### v1

@@ 4 Emergent Intelligence
Complex organisation in nature rarely arises from isolated units. From chemical reaction networks to neural circuits and ecosystems, intelligence is expressed through the collective behaviour of many interacting dynamical subsystems. Each component obeys simple local laws, yet the cooperation of many such components produces coherent structures, predictive control, and adaptive behaviour that appear qualitatively new. This phenomenon has long been recognised in both philosophy and the study of dynamical systems, from cellular automata such as Conway’s Game of Life [9] to coupled oscillators [17] and reaction–diffusion systems [28]. Despite this ubiquity, emergence has lacked a quantitative definition that links it directly to the underlying physics of information flow and conserved quantities.
In this section we provide such a definition. We extend the agent–environment framework to composite systems of many coupled subcomponents. By modelling the joint evolution of these subsystems and their shared environment through measurable transition kernels, we obtain a unified description of how coordination and coupling can increase the amount of useful work extracted per unit of irreversible information processed. This yields a physical and information-theoretic definition of emergent intelligence as the enhancement of goal-directed work that arises when interacting components collectively realise more goal-aligned work per nat of irreversible computation than they could in isolation. The following subsections formalise this framework and develop the corresponding metrics and bounds.
4.1 Multi-Component Agent-Environment Dynamics Model
@@
Let $(X^{(i)},\mathcal{X}^{(i)})$ for $i=1,\dots,N$ denote measurable state spaces of $N$ internal subsystems, and let $(E,\mathcal{E})$ denote the measurable state space of the external environment. The joint internal state is
$X:=X^{(1)}\!\times\!\cdots\!\times\!X^{(N)},\qquad\mathcal{X}:=\mathcal{X}^{(1)}\!\otimes\!\cdots\!\mathcal{X}^{(N)}.$ (45)
@@
$\Gamma_{X,i}:X^{(i)}\to\mathscr{X}^{(i)},\qquad\Gamma_{X}(x):=\big(\Gamma_{X,1}(x^{(1)}),\dots,\Gamma_{X,N}(x^{(N)})\big),$ (46)
and the environment exposes a port $\Gamma_{E}:E\to\mathscr{E}$. The pair $(\Gamma_{X}(X_{t}),\Gamma_{E}(E_{t}))$ forms the interface through which conserved-quantity exchange and information flow occur.
@@
$X_{t+1}^{(i)}$ $\sim K_{X^{(i)}}\!\left(\cdot\,\middle|\,\Gamma_{E}^{\mathcal{M}_{i}}(E_{t}),\,\Gamma_{X}^{\mathcal{N}_{i}}(X_{t})\right),$ (47)
$E_{t+1}$ $\sim K_{E}\!\left(\cdot\,\middle|\,E_{t},\Gamma_{X}(X_{t})\right),$ (48)
with conditional independences ensuring that all coupling occurs through the designated ports.
@@
$K(\mathrm{d}x^{\prime},\mathrm{d}e^{\prime}\mid x,e)$ $=\Big[\prod_{i=1}^{N}K_{X^{(i)}}(\mathrm{d}x^{(i)\prime}\mid\Gamma_{E}^{\mathcal{M}_{i}}(e),\Gamma_{X}^{\mathcal{N}_{i}}(x))\Big]K_{E}(\mathrm{d}e^{\prime}\mid e,\Gamma_{X}(x)),$ (49)
@@
The interface law is $\bar{\mu}_{t}:=(\Gamma_{X},\Gamma_{E})_{\#}\mu_{t}$.
A separable ensemble is obtained by removing inter-component port dependences in (47), yielding
$K_{X^{(i)}}^{\mathrm{sep}}(\mathrm{d}x^{(i)\prime}\mid\Gamma_{E}^{\mathcal{M}_{i}}(e)),$ (51)
and joint kernel
$K^{\mathrm{sep}}=\Big[\prod_{i=1}^{N}K_{X^{(i)}}^{\mathrm{sep}}\Big]\,K_{E}.$ (52)
The corresponding measure recursion defines $\mu_{t}^{\mathrm{sep}}$ and interface law $\bar{\mu}_{t}^{\mathrm{sep}}$, which serve as the uncoupled baseline.
4.2 Quantification of Emergence
The quantitative metrics introduced in §3—information capacity $\mathcal{C}_{T}$, identifiability capacity $\mathcal{F}_{T}$, and intelligence $\chi_{T}$—extend naturally to composite systems and retain their definitions when evaluated with respect to the joint law of the coupled ensemble. Their separable counterparts are computed analogously under the baseline dynamics. Coupling among subsystems alters both the total useful work performed and the rate at which irreversible information is processed in the conserved-quantity channel. The corresponding advantage over the separable baseline is
$\Delta\chi_{T}$ $=\chi_{T}-\chi_{T}^{\mathrm{sep}},$ (53)
with $\Delta\chi_{T}>0$ indicating emergent intelligence: the coupled system converts irreversible information into goal-directed work more efficiently than the same components operating independently.
To summarise this effect, define the dimensionless emergence index
$\mathcal{E}_{T}:=\frac{\Delta\chi_{T}}{\chi_{T}^{\mathrm{sep}}},\qquad\mathcal{E}_{T}>0\;\Longleftrightarrow\;\text{emergence}.$ (54)
The magnitude of $\mathcal{E}_{T}$ quantifies how strongly coupling enhances the system’s ability to translate irreversible information into useful work. Emergent organisation corresponds to coupling structures that improve work extraction per nat of irreversible computation, whether through energetic synergies (altering work pathways) or informational synergies (increasing predictive correlation across components). In this sense, emergent intelligence reflects the same physical principles governing single-agent intelligence, now amplified through cooperative structure and shared information flow.
5 Intelligence and Entropy
@@
Traditional discussions treat intelligence and entropy as opposing forces: intelligence as organisation and purpose, entropy as disorder and decay. Within the physical framework developed here, these concepts instead form a unified structure. Intelligence measures the efficiency with which a system converts irreversible information processing into goal-directed work, while entropy production quantifies the dissipative component of that information flow. The GENERIC decomposition introduced in §2.4 provides the natural setting for relating these quantities. Related connections between information processing, entropy production, and physical efficiency have been explored in nonequilibrium thermodynamics and the thermodynamics of computation [18, 6, 21].
5.1 Intelligence as Structured Reversible Flow with Selective Dissipation
@@
The internal dynamics of an intelligent agent admit a reversible-dissipative decomposition of the form
$\dot{x}_{t}=J(x_{t},t)\,\nabla H(x_{t},t)-R(x_{t},t)\,\nabla\Xi(x_{t},t),$ (55)
@@
where the reversible flow $J\nabla H$ preserves internal informational structure, while the dissipative flow $R\nabla\Xi$ induces irreversible contraction associated with conserved-quantity dissipation. This decomposition makes explicit the physical distinction between information-preserving internal dynamics and the irreversible processes required to exert control and influence over the environment.
The dissipative component generates a generalized entropy-production rate in the conserved-quantity channel,
$\dot{S}_{\mathrm{phys}}(t)=\langle\nabla\Xi(x_{t},t),\,R(x_{t},t)\,\nabla\Xi(x_{t},t)\rangle\;\geq\;0,$ (56)
with corresponding irreversible-information rate
$\dot{I}_{\mathrm{irr}}(t)=\dot{S}_{\mathrm{phys}}(t)/\alpha,$ (57)
where $\alpha$ denotes the fluctuation scale associated with the relevant conserved quantity. This irreversible-information flow represents the physical cost of interventions that break time-reversal symmetry and enable causal influence.
Goal-directed work arises from the interaction between internal dynamics and the environment and decomposes as
$\dot{W}_{\mathrm{goal}}(t)=-\langle\nabla G,\,J\nabla H\rangle+\langle\nabla G,\,R\nabla\Xi\rangle-\partial_{t}G.$ (58)
The reversible flow $J\nabla H$ enables the formation and maintenance of internal structure—such as memory, coordination, and predictive organization—that supports sustained control. The dissipative flow $R\nabla\Xi$, while costly, provides the irreversible interventions required to select actions, trigger environmental responses, and unlock external free energy along goal-aligned degrees of freedom.
Defining the instantaneous intelligence rate as
$\chi(t):=\frac{\dot{W}_{\mathrm{goal}}(t)}{\dot{I}_{\mathrm{irr}}(t)},$ (59)
we obtain a structural interpretation of intelligence: high intelligence corresponds to regimes in which irreversible dissipation is used sparingly and selectively to enable large amounts of goal-directed work mediated by the environment, while internal reversible dynamics preserve and reuse informational structure. Intelligence therefore increases not by eliminating dissipation, but by concentrating irreversible information processing into minimal interventions that exert maximal leverage over external entropy gradients. Such organisation is favoured because, for a given level of goal-directed work, it minimises the required rate of irreversible information processing while preserving reusable internal structure.
5.2 Conjecture: Intelligence as Constrained Entropy-Production Maximisation

### v2

@@ 6 Emergence
Complex organisation in nature rarely arises from isolated units. From chemical reaction networks to neural circuits and ecosystems, intelligent behaviour is expressed through the collective dynamics of many interacting subsystems. Each component obeys relatively simple local laws, yet their interaction gives rise to coherent structure, predictive control, and adaptive behaviour that are not attributable to any component in isolation. While such phenomena are ubiquitous across biological systems, a precise characterization of emergence grounded in the physics of information processing and conserved quantities has remained elusive.
In this section we provide such a characterization. Building on the framework developed in the preceding sections, we extend the agent–environment formalism to composite systems composed of many coupled subsystems. Emergence is treated here not as a new primitive, but as a macroscopic physical property: how coupling alters the thermodynamic efficiency with which irreversible and reversible computation are deployed, and in some cases, how the admissible system boundary itself is dynamically redefined to maximise this efficiency. We begin by giving a quantitative definition of emergent intelligence and then identify distinct emergent regimes corresponding to different patterns of change in macroscopic goal-directed work, preserved CCE structure, and system boundary geometry.
6.1 Multi-Component Agent–Environment Dynamics Model
@@
Let $(X^{(i)},\mathscr{X}^{(i)})$ for $i=1,\dots,N$ denote measurable state spaces of $N$ internal subsystems, and let $(E,\mathcal{E})$ denote the measurable state space of the external environment. The joint internal state is
$X:=X^{(1)}\!\times\!\cdots\!\times\!X^{(N)},\qquad\mathscr{X}:=\mathscr{X}^{(1)}\!\otimes\!\cdots\!\otimes\mathscr{X}^{(N)}.$ (94)
@@
$\Pi_{X,i}:X^{(i)}\to\Xi_{X},\qquad\Pi_{X}(x):=\big(\Pi_{X,1}(x^{(1)}),\dots,\Pi_{X,N}(x^{(N)})\big),$ (95)
and the environment exposes a port $\Pi_{E}:E\to\Xi_{E}$. The pair $(\Pi_{X}(X_{t}),\Pi_{E}(E_{t}))$ forms the physical interface through which conserved-quantity exchange and irreversible information flow occur.
@@
$X_{t+1}^{(i)}$ $\sim K_{X^{(i)}}\!\left(\cdot\,\middle|\,\Pi_{E}^{\mathcal{M}_{i}}(E_{t}),\,\Pi_{X}^{\mathcal{N}_{i}}(X_{t})\right),$ (96)
$E_{t+1}$ $\sim K_{E}\!\left(\cdot\,\middle|\,E_{t},\Pi_{X}(X_{t})\right),$ (97)
with conditional independences ensuring that all coupling occurs strictly through the designated ports.
@@
$K(\mathrm{d}x^{\prime},\mathrm{d}e^{\prime}\mid x,e)=\Big[\prod_{i=1}^{N}K_{X^{(i)}}(\mathrm{d}x^{(i)\prime}\mid\Pi_{E}^{\mathcal{M}_{i}}(e),\Pi_{X}^{\mathcal{N}_{i}}(x))\Big]K_{E}(\mathrm{d}e^{\prime}\mid e,\Pi_{X}(x)),$ (98)
@@
The interface law is $\bar{\mu}_{t}:=(\Pi_{X},\Pi_{E})_{\#}\mu_{t}$.
6.1.1 The Separable Baseline via Zero-Flux Boundary Conditions
To rigorously quantify emergence, we must define a separable baseline that isolates the subsystems from one another without altering their individual relationships with the external environment. Analogous to the passive boundary condition used to isolate causal footprint in §2.2, we define the separable ensemble by imposing a strict zero-flux boundary condition across all inter-component ports ($\mathcal{N}_{i}$).
Physically, this severs the exchange of conserved quantities and irreversible information between subsystems. Within the metriplectic formulation, this corresponds to setting all cross-system components of the Poisson and dissipative tensors to zero ($J_{ij}=0,R_{ij}=0$ for $i\neq j$).
The coupled joint kernel is therefore replaced by the completely separable kernel $K^{\mathrm{sep}}$:
$K^{\mathrm{sep}}(\mathrm{d}x^{\prime},\mathrm{d}e^{\prime}\mid x,e)=\Big[\prod_{i=1}^{N}K_{X^{(i)}}(\mathrm{d}x^{(i)\prime}\mid\Pi_{E}^{\mathcal{M}_{i}}(e))\Big]K_{E}(\mathrm{d}e^{\prime}\mid e,\Pi_{X}(x)),$ (100)
where $K_{X^{(i)}}$ now operates under the strictly enforced boundary condition that no active flux enters from any internal neighbor $j\in\mathcal{N}_{i}$. The resulting measure $\mu_{t}^{\mathrm{sep}}$ defines the uncoupled reference trajectory. Any deviation from this baseline—manifesting as an increase in efficiency ($\Delta\chi_{T}$) or preserved CCE structure ($\Delta\kappa_{T}$)—is therefore strictly attributable to the macroscopic physical coupling of the subsystems.
6.2 Quantitative Definition of Emergent Intelligence and Consciousness
Within the composite setting above, causal intelligence and consciousness are evaluated as defined in §3.1 and §3.2 using the joint law of the coupled ensemble. To isolate the physical contribution of coupling, we compare the coupled system to the corresponding separable baseline.
Crucially, because intelligence is defined as a ratio of integrated fluxes rather than the time integral of an instantaneous ratio, emergent contributions over a finite horizon $[0,T]$ must be computed directly from the difference of the cumulative causal efficiencies:
$\Delta\chi_{T}:=\chi_{T}-\chi_{T}^{\mathrm{sep}}=\frac{W_{\mathrm{causal},T}}{I_{\mathrm{irr},T}}-\frac{W_{\mathrm{causal},T}^{\mathrm{sep}}}{I_{\mathrm{irr},T}^{\mathrm{sep}}}$ (101)
$\Delta\kappa_{T}:=\kappa_{T}-\kappa_{T}^{\mathrm{sep}}=\frac{W_{\mathrm{causal},T}}{I_{\mathrm{rev},T}}-\frac{W_{\mathrm{causal},T}^{\mathrm{sep}}}{I_{\mathrm{rev},T}^{\mathrm{sep}}}$ (102)
Thus, $\Delta\chi_{T}>0$ indicates net emergent causal intelligence over $[0,T]$, and $\Delta\kappa_{T}>0$ indicates net emergent causal consciousness over $[0,T]$.
Goal-Directed Emergence
When evaluating emergence with respect to a specific task, we apply the goal-directed projection established previously to yield the functional emergent metrics:
$\Delta\chi_{\mathrm{goal},T}:=\chi_{\mathrm{goal},T}-\chi_{\mathrm{goal},T}^{\mathrm{sep}}=\frac{W_{\mathrm{goal},T}}{I_{\mathrm{irr},T}}-\frac{W_{\mathrm{goal},T}^{\mathrm{sep}}}{I_{\mathrm{irr},T}^{\mathrm{sep}}}$ (103)
$\Delta\kappa_{\mathrm{goal},T}:=\kappa_{\mathrm{goal},T}-\kappa_{\mathrm{goal},T}^{\mathrm{sep}}=\frac{W_{\mathrm{goal},T}}{I_{\mathrm{rev},T}}-\frac{W_{\mathrm{goal},T}^{\mathrm{sep}}}{I_{\mathrm{rev},T}^{\mathrm{sep}}}$ (104)
Throughout this section, all quantities are understood to be evaluated relative to the same admissible system boundary and operational coarse-graining, so that changes reflect genuine efficiency gains from coupling rather than a change of description.
For convenience, we also introduce the dimensionless horizon-$T$ indices:
$\mathcal{E}_{T}:=\frac{\Delta\chi_{T}}{\chi_{T}^{\mathrm{sep}}},\quad\mathcal{C}_{T}:=\frac{\Delta\kappa_{T}}{\kappa_{T}^{\mathrm{sep}}}.$ (105)
Their signs distinguish the types of emergence:
$\mathcal{E}_{T}>0$ $\iff\text{emergent intelligence},$ (106)
$\mathcal{C}_{T}>0$ $\iff\text{emergent consciousness}.$ (107)
These quantities measure the relative strength of the emergent effects and allow comparison across physically heterogeneous systems and scales.
6.3 Characteristic Regimes of Emergent Intelligence
Emergent phenomena may be classified according to which of the quantities $\chi$, $\kappa$, and the boundary $\Pi_{X}(t)$ change under coupling, and on what timescale. Three characteristic regimes are distinguished.
For the regime classification below, we compare the coupled system to its separable baseline by evaluating the sign and magnitude of the cumulative emergent efficiencies $\Delta\chi_{T}$ and $\Delta\kappa_{T}$ over the horizon $[0,T]$, alongside the dynamical evolution of the admissible boundary $\Pi_{X}(t)$.
6.3.1 Fast emergent coordination
In this regime, the admissible system boundary $\Pi_{X}$ remains fixed and the amount of preserved internal structure does not increase significantly over the horizon considered. Emergence manifests as a rapid increase in intelligence due to more effective real-time routing of irreversible information processing:
$\Delta\chi_{T}>0\,\qquad\Delta\kappa_{T}\approx 0,\qquad\Pi_{X}(t)=\Pi_{X,0}.$ (108)
Such events occur on short timescales and correspond to sudden coordination, synchronisation, or acute adaptive responses. In neural systems, this includes reflexive learning and extreme-stimulus avoidance, where intelligence increases transiently by pooling irreversible control interventions without immediate consolidation of long-term structure.
6.3.2 Slow structural consolidation
Here the system boundary remains fixed, but coupling leads to a gradual increase in preserved internal structure. The coupled subsystems collectively establish metastable CCE basins that would be thermodynamically inaccessible to individual components. Over long horizons, this reduces redundant irreversible computation and thereby increases intelligence:
$\Delta\chi_{T}>0\,\qquad\Delta\kappa_{T}>0,\qquad\Pi_{X}(t)=\Pi_{X,0}.$ (109)
This regime corresponds to memory formation, skill acquisition, and long-term learning, in which the generalised Landauer cost is invested once to construct stable structure that supports cheaper goal-directed behaviour thereafter.
6.3.3 Boundary-changing emergence
In this regime, coupling triggers the boundary optimization process described in §2.4, leading to a dynamic redefinition of the admissible system boundary itself:
$\Pi_{X}(t):\ \Pi_{X,0}\;\longrightarrow\;\Pi_{X}^{\prime}(t),$ (110)
giving rise to a new, integrated macroscopic system at a higher level of organisation. Over a horizon $[0,T]$, this corresponds to a time-dependent admissible boundary $\Pi_{X}(t)$. The integrated emergent contributions are therefore
$\Delta\chi_{T}>0,\qquad\Delta\kappa_{T}>0,$ (111)
and must be interpreted relative to this evolving description.
Such events include incorporative processes such as endosymbiosis, as well as recombinative processes such as sexual reproduction, in which information from multiple systems is combined to instantiate a new dynamically closed system. In these cases, emergence is not merely an efficiency gain within a fixed system, but a physical transition in system individuation—a shift in the geometric boundary of the agent itself to maximise intelligence.
Together, these regimes demonstrate that emergence within the present framework is neither monolithic nor mysterious. It arises through well-defined changes in efficiency, preserved CCE structure, and optimal system boundaries, governed strictly by the same physical principles that underlie single-system intelligence.
7 Intelligence and Generalised Dissipation: The Physical Drive for Cognition
@@
Traditional discussions treat intelligence and physical decay as opposing forces: intelligence as organisation and purpose, dissipation as disorder and breakdown. Within the physical framework developed here, these concepts instead form a unified structure. Intelligence measures the efficiency with which a system converts irreversible information processing into goal-directed work, while generalised dissipation quantifies the strict physical cost of that computation under conservation laws.
To understand why intelligent systems emerge and evolve in the physical universe, we must look beyond classical thermodynamics to the generalised dynamics of open, far-from-equilibrium systems constrained by the limits of physical encoding. Related connections between information processing, entropy production, and physical efficiency have been explored in nonequilibrium thermodynamics and the thermodynamics of computation [15, 5, 17].
7.1 Intelligence as Structured Reversible Flow with Selective CCE Collapse
@@
The internal dynamics of an intelligent agent admit a reversible-dissipative decomposition of the form:
$\dot{x}_{t}=J(x_{t},t)\,\nabla H(x_{t},t)-\mathcal{R}(x_{t},t)\,\nabla\Xi(x_{t},t)$ (112)
@@
where the reversible flow $J\nabla H$ preserves internal informational structure, while the dissipative flow $\mathcal{R}\nabla\Xi$ induces irreversible contraction associated with conserved-quantity dissipation. This decomposition makes explicit the physical distinction between information-preserving internal dynamics and the irreversible processes required to exert control and influence over the environment.The dissipative component generates a generalised dissipation rate within the specific conserved-quantity channel stabilising the system:
$\dot{S}_{\mathrm{phys}}(t)=\langle\nabla\Xi(x_{t},t),\,\mathcal{R}(x_{t},t)\,\nabla\Xi(x_{t},t)\rangle\geq 0$ (113)
with a corresponding irreversible-information rate:
$\dot{I}_{\mathrm{irr}}(t)=\dot{S}_{\mathrm{phys}}(t)/\alpha$ (114)
where $\alpha$ denotes the fluctuation scale associated with the relevant conserved quantity (e.g., $k_{B}$ for thermal baths, $\hbar$ for quantum mechanical channels). This irreversible-information flow represents the strict physical exhaust of interventions that break time-reversal symmetry to collapse a Conservation-Congruent Encoding (CCE).The Causal Power DecompositionMacroscopic causal work arises from the interaction between the agent’s internal dynamics and the environment via the boundary fluxes. Let $\mathcal{F}_{\mathrm{causal}}:=\nabla_{x}{W}_{\mathrm{causal}}$ denote the effective causal gradient, representing the sensitivity of the instantaneous macroscopic causal power to changes in the agent’s internal physical state. The active contribution of the agent’s dynamics to the causal work rate decomposes linearly over the internal metriplectic flow:
$\dot{W}_{\mathrm{causal}}(t)=\langle\mathcal{F}_{\mathrm{causal}},\,J(x_{t},t)\nabla H(x_{t},t)\rangle-\langle\mathcal{F}_{\mathrm{causal}},\,\mathcal{R}(x_{t},t)\nabla\Xi(x_{t},t)\rangle$ (115)
The first term, $\langle\mathcal{F}_{\mathrm{causal}},\,J\nabla H\rangle$, represents the causal power enabled by reversible, structure-preserving dynamics. This flow allows the formation and maintenance of internal CCEs—such as memory, coordination, and predictive organization—that route conserved quantities and support sustained control without producing generalised dissipation.Conversely, the second term, $\langle\mathcal{F}_{\mathrm{causal}},\,\mathcal{R}\nabla\Xi\rangle$, represents the causal power unlocked via irreversible interventions. While physically costly, this dissipative flow provides the symmetry-breaking steps required to collapse encodings, select actions, and unlock external gradients along causal degrees of freedom.Defining the instantaneous causal intelligence rate as:
$\chi(t):=\frac{\dot{W}_{\mathrm{causal}}(t)}{\dot{I}_{\mathrm{irr}}(t)}$ (116)
we obtain a structural interpretation of baseline intelligence. High intelligence corresponds to regimes in which the irreversible collapse of CCEs is used sparingly and selectively to enable large amounts of causal work mediated by the environment.The Goal-Directed Power DecompositionWhen evaluating task-specific efficiency, we project this dynamic influence onto the goal-aligned subspace. Let $\mathcal{F}_{\mathrm{goal}}:=\nabla_{x}{W}_{\mathrm{goal}}$ denote the effective task gradient. The active contribution to the goal-directed work rate follows the identical geometric decomposition:
$\dot{W}_{\mathrm{goal}}(t)=\langle\mathcal{F}_{\mathrm{goal}},\,J(x_{t},t)\nabla H(x_{t},t)\rangle-\langle\mathcal{F}_{\mathrm{goal}},\,\mathcal{R}(x_{t},t)\nabla\Xi(x_{t},t)\rangle$ (117)
yielding the instantaneous goal-directed intelligence rate:
$\chi_{\mathrm{goal}}(t):=\frac{\dot{W}_{\mathrm{goal}}(t)}{\dot{I}_{\mathrm{irr}}(t)}$ (118)
Under both causal and goal-directed evaluations, intelligence increases not by eliminating dissipation, but by concentrating the generalised Landauer cost into minimal interventions that exert maximal leverage over external structural gradients.
7.2 Intelligence as Constrained Throughput Maximization